a rectangular field is twice as long as it is wide, if the perimeter of the field is 96 yards, what are the dimensions ?
Answer by oberobic(2304)

P = perimeter
L = length
W = width
.
P = 96
.
P = 2(L+W)
L+W = 96/2 = 48
.
L = 2W
.
substitute for L
.
2W + W = 48
.
3W = 48
.
W = 16
.
substitute W
.
L = 48-W = 32
.
L is twice W
.
Answer: Length = 32 and width = 16.
.
Done.